# CoTAR: Chain-of-Thought Attribution Reasoning with Multi-level Granularity

### v1

@@ Abstract
State-of-the-art performance in QA tasks is currently achieved by systems employing Large Language Models (LLMs), however these models tend to hallucinate information in their responses. One approach focuses on enhancing the generation process by incorporating attribution from the given input to the output. However, the challenge of identifying appropriate attributions and verifying their accuracy against a source is a complex task that requires significant improvements in assessing such systems. We introduce an attribution-oriented Chain-of-Thought reasoning method to enhance the accuracy of attributions. This approach focuses the reasoning process on generating an attribution-centric output. Evaluations on two context-enhanced question-answering datasets using GPT-4 demonstrate improved accuracy and correctness of attributions. In addition, the combination of our method with finetuning enhances the response and attribution accuracy of two smaller LLMs, showing their potential to outperform GPT-4 in some cases.
@@ 1 Introduction
Figure 1: Usage of CoT for attribution-based answers.
@@ 2.2 Chain-of-Thought Attribution Reasoning
Level CoT Answer Quality Citation Quality BERT HEM RL Avg. ALCE F1 CSCA DOC F1 SEM-F1 Avg. Span None 91.71 67.58 66.08 75.12 77.06 81.15 85.33 67.57 77.78 Span 92.06 66.95 68.55 75.85 78.83 89.88 86.54 72.23 81.87 Sent. 91.87 66.88 68.76 75.84 84.76 69.69 86.38 70.52 77.84 Pass. 91.91 66.89 67.67 75.49 79.21 86.94 86.49 69.87 80.63 Sent. None 91.47 62.48 62.18 72.04 82.79 - 85.32 59.20 75.77 Span 91.92 62.06 64.76 72.91 82.39 - 86.71 61.74 76.94 Sent. 91.96 64.59 67.36 74.64 86.42 - 86.51 64.44 79.12 Pass. 91.69 61.19 64.42 72.43 84.25 - 85.92 61.79 77.32 Pass. None 91.55 61.13 62.76 71.81 75.38 - 85.19 44.31 68.29 Span 91.90 61.66 64.70 72.75 71.60 - 86.17 43.41 67.06 Sent. 91.84 62.43 65.32 73.20 70.63 - 86.41 42.59 66.54 Pass. 91.69 60.40 63.88 71.99 73.28 - 85.64 44.12 67.68
@@ 3 Evaluation Metrics
For answer quality , we use the n-grams based method ROUGE-L (RL), and the semantic method BERTScore (BERT) Zhang* et al. (2020). For hallucination evaluation, we use the HEM111vectara/hallucination_evaluation_model model, which was finetuned on NLI datasets, and predict whether two texts are factually consistent. For citation quality, we measure the quality of the cited content; we propose specific metrics for each of the citation levels. We note that span citations must match the source passages precisely, while sentence and passage attributions are not required to do so. We use the SEM-F1 metric proposed by Schuster et al. (2023) for n-gram token level similarity between the cited content. In addition, we include the Citation Precision/Recall introduced in Gao et al. (2023), and combine them into ALCE F1, by using the harmonic mean. We also measure the F1 score over passage indices between the cited passages and the passages cited by the expected answer, denoted as DOC F1, and the Correct Span Citation Attribution (CSCA) which indicates whether a predicted span is a direct span from the attributed passage.
@@ 4.4 Method Comparison
5 Experimental Results
To examine how CoT affects the model, we inspect the GPT-4 case, and provide a detailed overview of the metrics in Table 1. In the answer metrics, the differences between the various CoT methods and the standard run (shown as "None") are not significant, aside from the sentence level, where the sentence CoT method produces better scores across all metrics. For the citation quality metrics, the span and sentence level produce better scores than their counterparts when using the appropriate CoT methods. In particular, the CSCA metric in the span level case showcases a significant advantage, mainly since the model is guided to produce the spans relevant for the answer in advance, thus allowing the model to accurately reference them. At the passage level, the SEM-F1 and ALCE-F1 scores showcase an advantage for the standard run. We hypothesize that the addition of CoT for such a broad citation task is less helpful than in the more granular sentence and span level citations. We observe that the GPT-4 model performed best in the span level when using the span CoT method, and the same applies for the sentence level with sentence CoT. In the passage level, while the CoT methods improve the answer quality, they do not aid in the citation quality.
We showcase similar results for all the models, across the various citation levels and CoT methods in Tables 7 and 8. For each setting, we display all metrics and category averages for answer and citation quality, respectively.
@@ 5 Experimental Results
Table 2: Average answer and citation quality results, per citation level, per CoT method, for QuoteSUM and MSMARCO. Rows: Citation levels. Columns: CoT method used and metrics averages per level.
@@ 6 Related Work
Several recent studies have suggested different techniques for text generation including citation attribution, with varying degrees of detail. Attributing relevant information to web pages was done in Thoppilan et al. (2022); Liu et al. (2023). Menick et al. (2022) attributes the response to specific snippets from the source passages. In particular, Gao et al. (2023) assigns attribution for each produced sentence to one or several input passages. Our sentence level citation approach and the evaluation metrics are based on this work.
@@ Appendix C Additional Results
1:Input: Answer $a$, passages $\{p_{i}\}$
9:for $sc\in SC(a)$ do
10:     if  $!\text{IsMarked}(sc,a,a_{\text{marked}})$ then
11:         $\text{Mark}(sc,a,a_{\text{marked}})$
12:     end if
13:end for
15:return $a_{\text{span}}$

### v2

@@ Abstract
State-of-the-art performance in QA tasks is currently achieved by systems employing Large Language Models (LLMs), however these models tend to hallucinate information in their responses. One approach focuses on enhancing the generation process by incorporating attribution from the given input to the output. However, the challenge of identifying appropriate attributions and verifying their accuracy against a source is a complex task that requires significant improvements in assessing such systems. We introduce an attribution-oriented Chain-of-Thought reasoning method to enhance the accuracy of attributions. This approach focuses the reasoning process on generating an attribution-centric output. Evaluations on two context-enhanced question-answering datasets using GPT-4 demonstrate improved accuracy and correctness of attributions. In addition, the combination of our method with finetuning enhances the response and attribution accuracy of two smaller LLMs, showing their potential to outperform GPT-4 in some cases. [^1] Our code is publicly available for reproduction: https://github.com/mosheber/cotar.git
@@ 1 Introduction
Figure 1: Usage of CoT for attribution-based answers. We either instruct the model, using fewshot examples, or finetune the model, to produce a detailed list of the salient information from each passage. Each entry can be either on the passage, sentence, and even span level. Finally, the model produces a coherent and faithful answer.
@@ 2.2 Chain-of-Thought Attribution Reasoning
Level CoT Answer Quality Citation Quality
BERT HEM RL Avg. ALCE F1 CSCA DOC F1 SEM-F1 Avg.
Span None 91.71 67.58 66.08 75.12 77.06 81.15 85.33 67.57 77.78
Span 92.06 66.95 68.55 75.85 78.83 89.88 86.54 72.23 81.87
Sent. 91.87 66.88 68.76 75.84 84.76 69.69 86.38 70.52 77.84
Pass. 91.91 66.89 67.67 75.49 79.21 86.94 86.49 69.87 80.63
Sent. None 91.47 62.48 62.18 72.04 82.79 - 85.32 59.20 75.77
Span 91.92 62.06 64.76 72.91 82.39 - 86.71 61.74 76.94
Sent. 91.96 64.59 67.36 74.64 86.42 - 86.51 64.44 79.12
Pass. 91.69 61.19 64.42 72.43 84.25 - 85.92 61.79 77.32
Pass. None 91.55 61.13 62.76 71.81 75.38 - 85.19 44.31 68.29
Span 91.90 61.66 64.70 72.75 71.60 - 86.17 43.41 67.06
Sent. 91.84 62.43 65.32 73.20 70.63 - 86.41 42.59 66.54
Pass. 91.69 60.40 63.88 71.99 73.28 - 85.64 44.12 67.68
@@ 3 Evaluation Metrics
For answer quality , we use the n-grams based method ROUGE-L (RL), and the semantic method BERTScore (BERT) Zhang* et al. (2020). For hallucination evaluation, we use the HEM222vectara/hallucination_evaluation_model model, which was finetuned on NLI datasets, and predict whether two texts are factually consistent. For citation quality, we measure the quality of the cited content; we propose specific metrics for each of the citation levels. We note that span citations must match the source passages precisely, while sentence and passage attributions are not required to do so. We use the SEM-F1 metric proposed by Schuster et al. (2023) for n-gram token level similarity between the cited content. In addition, we include the Citation Precision/Recall introduced in Gao et al. (2023), and combine them into ALCE F1, by using the harmonic mean. We also measure the F1 score over passage indices between the cited passages and the passages cited by the expected answer, denoted as DOC F1, and the Correct Span Citation Attribution (CSCA) which indicates whether a predicted span is a direct span from the attributed passage.
@@ 4 Experimental Setup
In order to properly manage our experimental setup, we utilized RAGFoundry, as introduced in Fleischer et al. (2024). This allowed us to create the data preprocessing flows, train our models, conduct inference, and evaluate on the various metrics in a customizable and seamless manner. Our codebase provides a detailed description of how to reproduce the results using the custom steps we have incorporated with the framework.
@@ 4.3 Models
1:Input: Answer $a$, passages $\{p_{i}\}$
9:for $sc\in SC(a)$ do
10:     if  $!\text{IsMarked}(sc,a,a_{\text{marked}})$ then
11:         $\text{Mark}(sc,a,a_{\text{marked}})$
12:     end if
13:end for
15:return $a_{\text{span}}$
@@ 4.4 Method Comparison
Table 2: Average answer and citation quality results, per citation level, per CoT method, for QuoteSUM and MSMARCO. Columns: Citation levels. Rows: CoT method used and metrics averages per level.
Idx. Expected Answer No-CoT Span-CoT
75
(2) [ 2 general paralysis
of the insane (GPI) ]
GPI stands for [ 2 general
paralysis of the
insane ], [ 3 Georgia Budget
and Policy Institute ], and [ 5
Global Residence Program Index ].
Lets analyze the relevant spans
of information from the input
passages.
From passage [2], the relevant
spans are the following:
* general paralysis of the insane
Thus, the final answer is:
GPI stands for [ 2 general paralysis
of the insane ].
941
Some sources state that
[ 1 "Born on the Fourth
of July" ] lasts [ 1 for
two hours and
twenty-five minutes. ]
[ 2 "Born on the Fourth of July"
was written in Santa Monica, California,
during the fall of 1974
in exactly one month, three
weeks and two days. ]
[ 1 Born on the Fourth of July ]
is the title of [ 1 Stone’s movie ],
which runs [ 1 two hours and
twenty-five minutes ].
[ 2 "Born on the Fourth of July"
was written in Santa Monica,
California, during the fall of 1974
in exactly one month, three
weeks and two days. ]
Table 3: Examples of answer results when utilizing the span CoT method, as opposed to not using CoT at all. The examples are taken from the QuoteSUM test set, using the finetuned Mistral-7B models in both cases.
5 Experimental Results
To examine how CoT affects the model, we inspect the GPT-4 case, and provide a detailed overview of the metrics in Table 1. In the answer metrics, the differences between the various CoT methods and the standard run (shown as "None") are not significant, aside from the sentence level, where the sentence CoT method produces better scores across all metrics. For the citation quality metrics, the span and sentence level produce better scores than their counterparts when using the appropriate CoT methods.
In particular, the CSCA metric in the span level case showcases a significant advantage, mainly since the model is guided to produce the spans relevant for the answer in advance, thus allowing the model to accurately reference them.
In Table 3, we showcase some examples of how the span CoT method allows the model to better quote spans from the original text. The first example shows that it allows the model to avoid quoting irrelevant spans. The second one showcases that while it does at times quote the wrong information, it still manages to quote the right information as well, as opposed to not using CoT at all.
These results also support our claims, as shown in the DOC F1 metric, showcasing a significant advantage for our method across the various citation levels. Since DOC F1 is lower for missing and excessive citations, our superior performance in it demonstrates how our approach improves upon those aspects. At the passage level, the SEM-F1 and ALCE-F1 scores showcase an advantage for the standard run. We hypothesize that the addition of CoT for such a broad citation task is less helpful than in the more granular sentence and span level citations. We observe that the GPT-4 model performed best in the span level when using the span CoT method, and the same applies for the sentence level with sentence CoT. In the passage level, while the CoT methods improve the answer quality, they do not aid in the citation quality.
We showcase similar results for all the models, across the various citation levels and CoT methods in Tables 8 and 9. For each setting, we display all metrics and category averages for answer and citation quality, respectively.
@@ 6 Related Work
Several recent studies have suggested different techniques for text generation including citation attribution, with varying degrees of detail. Attributing relevant information to web pages was done in Thoppilan et al. (2022); Liu et al. (2023); Bohnet et al. (2022). Menick et al. (2022) attributes the response to specific snippets from the source passages. In particular, Gao et al. (2023) assigns attribution for each produced sentence to one or several input passages. Our sentence level citation approach and the evaluation metrics are based on this work.